OpenStudy (anonymous):

Solve \[2e^{x}(e^{x}-e)=3e^{x}-4\]

OpenStudy (anonymous):

\[2e^{x}(e^{x}-3)=3e^{x}-4\]

OpenStudy (freckles):

distribute on the left hand side first

OpenStudy (freckles):

then put everything on one side you should see a quadratic in terms of e^x

OpenStudy (anonymous):

i really dont get this at all, like i got nothing :(

OpenStudy (anonymous):

Awesome. Thanks!

OpenStudy (anonymous):

for what?

OpenStudy (freckles):

Let me know if you need further help @Dscdago

OpenStudy (anonymous):

help mee pleasee

OpenStudy (anonymous):

Theorem: The segment joining the midpoints of two sides of a triangle is parallel to the third side and half its length. A two-column proof of the theorem is shown, but the statement and reasons are not in correct order. A triangle with vertices A is at 6, 8. B is at 2, 2. C is at 8, 4. Segment DE with point D on side AB and point E is on side BC Statement Reason I The coordinates of point D are (4, 5) and coordinates of point E are (5, 3) By the midpoint formula II Segment DE is half the length of segment AC. By substitution III Length of segment DE is Square root of 5 and length of segment AC is 2 multiplied by the square root of 5 By the distance formula IV Segment DE is parallel to segment AC. Slopes of parallel lines are equal. V Slope of segment DE is -2 and slope of segment AC is -2. By the slope formula Which is the most logical order of statements and reasons for the proof? II, III, I, IV, V I, III, II, V, IV I, III, II, IV, V III, II, I, V, IV

OpenStudy (anonymous):

@freckles

OpenStudy (anonymous):

@ dan815

OpenStudy (anonymous):

@dan815

OpenStudy (anonymous):

@freckles how does the quadratic look like? to make sure

OpenStudy (freckles):

\[2e^x(e^x-3)=3e^x-4 \\ 2e^{2x}-6e^x=3e^x-4 \\ 2e^{2x}-6e^{x}-3e^{x}+4=0 \\ 2e^{2x}-9e^{x}+4=0\]

OpenStudy (freckles):

where e^x=u

OpenStudy (anonymous):

Got it now. Thanks again!